OpenStudy (anonymous):

solve. 1. 5log(x-2)=11 2. 12=10^(x+5)-7

OpenStudy (anonymous):

\[12=10^{x+5}-7\] \[19=10^{x+5}\] \[\log19=x+5\] \[\log19 - 5=x\] \[-3.721\approx x\]

OpenStudy (anonymous):

5log(x-2)=11 log(x-2)=11/5 10^(11/5)=x-2 10^(11/5)+2 = x \[160 \approx x\]